Measure Maps Rizzo Goal Inc. produces and sells hockey equipment, often custom made for online orders....

  1. Measure Maps

    Rizzo Goal Inc. produces and sells hockey equipment, often custom made for online orders. The company has the following performance metrics on its balanced scorecard: days from ordered to delivered, number of shipping errors, customer retention rate, and market share. A measure map illustrates that the days from ordered to delivered and the number of shipping errors are both expected to directly affect the customer retention rate, which affects market share. Additional internal analysis finds that:

    • Every shipping error over 3 shipping errors per month reduces the customer retention rate by 1.5%.
    • On average, each day above 3 days from ordered to delivered yields a reduction in the customer retention rate of 1%.
    • Each day before 3 days from order to delivery yields an increase in the customer retention rate of 1%, on average.
    • Rizzo Goal Inc.’s current customer retention rate is 80%.
    • The company estimates that for every 1% increase or decrease in the customer retention rate, market share changes 0.5% in the same direction.
    • Rizzo Goal Inc.’s current market share is 22.4%.

    Ignoring any other factors, if the company has 6 shipping errors this month and an average of 3.5 days from ordered to delivered.

    a. Determine the new customer retention rate. Round your answer to one decimal place.
    %

    b. Determine the new market share that Rizzo Goal Inc. expects to have. Round your answer to one decimal place.
    %
Answer #1

The measure maps depict which metrics are leading and lagging indicators of performance. It shows how strategic objective contributes to overall mission or strategy.

a.

day above 3 days from order to delivery = 3.5 – 3 = 0.5 days

decrease in customer retention rate = 0.5* 1% = 0.5%

errors above three per month = 6 – 3 = 3

decrease in customer retention rate = 3 * 1.5% = 4.5%

New customer retention rate = 80% - 0.5% - 4.5% = 75%

b.

total decrease in customer retention rate = 0.5+4.5 = 5%

decrease in market share = 5% *0.5 = 2.5%

New market share = 22.4 – 2.5 % = 19.9%